OpenStudy (anonymous):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 3, – 25), and x intercepts at x = -8 and x = 2. (Do not include the negative sign in your answer.) y = x2 + 6x − ___

OpenStudy (anonymous):

the missing term there is going to be a constant. lets call it 'c' We are also given a point on the graph (-3,-25) We can plug this in and solve for c. y = x2 + 6x + c -25 = -3^2 + 6(-3) + c -25 = 9 + -18 + c -16 = c We can check it by using the x intercepts given. y will be zero at these points. 0 = 2^2 + 6(2) - 16 0 = 4 + 12 - 16 0=0 We are correct. Hope this helped

OpenStudy (anonymous):

quick note. I have a "+ c" where the equation in your question has a "- c" so the answer they are looking for would be 16